An insurance company is analyzing the following three bonds, each with five years to maturity, annual interest payments, and is using duration as the measure of interest rate risk. What is the duration of each of the three bonds? (Do not round intermediate calculations. Round your answers to 2 decimal places. (e.g., 32.16)) Duration of the bond $10,000 par value, coupon rate-9.5%, rb-0.15 |$10,000 par value, coupon rate 11.5%, rb 0.15 |$10,000 par value, coupon rate-13.5%, rb-0.15 4.01 years 3.91 3 82。 C.
Answer #1

a.

Period

Cash Flow from Bond

Discounting factor = 1/(1+R)^N

PV of the cash flows = Cash flow x Df

Weighted cash flow = Period x Cash flow

Present value of weighted cash flow = Weighted Cash flow x Df

N

CF

Df = 1/(1+0.15)^N

PV = CF x Df

WCF = CF x N

WPV = WCF x Df

1

950.00

0.8696

826.0870

950.0000

826.0870

2

950.00

0.7561

718.3365

1900.0000

1436.6730

3

950.00

0.6575

624.6404

2850.0000

1873.9213

4

950.00

0.5718

543.1656

3800.0000

2172.6623

5

10950.00

0.4972

5444.0853

54750.0000

27220.4263

Total = P = Price =

8156.3147

Total = Weighted Price = WP

33529.7698

Duration or Macaulay Duration = (WP/P) =

4.11 Years

b.

Period

Cash Flow from Bond

Discounting factor = 1/(1+R)^N

PV of the cash flows = Cash flow x Df

Weighted cash flow = Period x Cash flow

Present value of weighted cash flow = Weighted Cash flow x Df

N

CF

Df = 1/(1+0.15)^N

PV = CF x Df

WCF = CF x N

WPV = WCF x Df

1

1150.00

0.8696

1000.0000

1150.0000

1000.0000

2

1150.00

0.7561

869.5652

2300.0000

1739.1304

3

1150.00

0.6575

756.1437

3450.0000

2268.4310

4

1150.00

0.5718

657.5162

4600.0000

2630.0649

5

11150.00

0.4972

5543.5206

55750.0000

27717.6030

Total = P = Price =

8826.7457

Total = Weighted Price = WP

35355.2294

Duration or Macaulay Duration = (WP/P) =

4.01 Years

c.

Period

Cash Flow from Bond

Discounting factor = 1/(1+R)^N

PV of the cash flows = Cash flow x Df

Weighted cash flow = Period x Cash flow

Present value of weighted cash flow = Weighted Cash flow x Df

N

CF

Df = 1/(1+0.15)^N

PV = CF x Df

WCF = CF x N

WPV = WCF x Df

1

1350.00

0.8696

1173.9130

1350.0000

1173.9130

2

1350.00

0.7561

1020.7940

2700.0000

2041.5879

3

1350.00

0.6575

887.6469

4050.0000

2662.9407

4

1350.00

0.5718

771.8669

5400.0000

3087.4675

5

11350.00

0.4972

5642.9559

56750.0000

28214.7797

Total = P = Price =

9497.1767

Total = Weighted Price = WP

37180.6889

Duration or Macaulay Duration = (WP/P) =

3.91 Years